You are on the prom committee. You and the other members are deciding how many student and adult tickets to sell. If the committee needs to raise $1,244 dollars and you have decided to sell student tickets for $2.00 each and adult tickets for $5.00. There are 421 people attending, how many student tickets have you sold.
Answer by stanbon(48546)

Let number of students tikets be "x"
The number of adult tickets will be "421-x"
Value of student tickets is 200x cents
Value of adult tickets is 500(421-x)=210500-500x cents
EQUATION:
Value of student + Value of adult tickets=124400 cents
200x + 210500-500x = 124400
-300x= -86100
x=287 (number of student tickets)
421-x=134 (number of adult tickes)
Cheers,
Stan H.